OpenStudy (anonymous):

plz help John deposited $20,000 for 2 years with a friend who gave him 12% simple interest. How much more interest will he receive if he deposits the same sum in a bank that offers 12% compound interest? A.$288 B.$290 C.$240 D.$200

OpenStudy (anonymous):

I'm pretty certain simple interest is like: $20,000 * 2 years * 12% = $4,800

OpenStudy (anonymous):

are you saying the answer is 4,800

OpenStudy (anonymous):

No... that's just what John gets from the simple interest account. You have to compare it to what he would get from compound interest to see the difference.. how much more he would make.

OpenStudy (anonymous):

so divide it by two

OpenStudy (anonymous):

Why? Simple interest just means the bank will pay him a percentage of his total balance for each year he has his money there. Compound interest means they pay him a percentage after each year, but next year, they calculate the interest based on the amount of money he has PLUS the interest they paid him last year... so each year, the interest payment grows because the account balance grows every time they pay interest. Check these two links out for some explanations.. Simple interest: http://math.about.com/od/businessmath/ss/Interest.htm Compound interest: http://math.about.com/library/weekly/aa042002a.htm

OpenStudy (anonymous):

is it B

OpenStudy (anonymous):

i got the answer its A.288 thanks anyway

OpenStudy (anonymous):

Simple interest: I = P * r * t where I is interest, P is principal ($20,000), r is rate (12%), and t is time (2 years) The only compound interest formula I found was M = P(1 + r)^t but that gives you the total with interest, so you have to subtract 20000 to get just the interest So simple interest = 20000 * .12 * 2 compound = 20,000*(1 + 0.12)^2 - 20,000

OpenStudy (anonymous):

yes, that's what I got too. I wasn't sure where he got "B" needed to double check what I had done to be sure I didn't miss something.